# SelfDefend: LLMs Can Defend Themselves against Jailbreaking in a Practical Manner

### v1

@@ 7.2 Defense Effectiveness
We now discuss the defense effectiveness of our proposed technique. In line with Table IV, we discuss the results over GPT-3.5, GPT-4, Llama-2-7b-chat, and a “double shadow stack” setting, respectively.
@@
Results on GPT-3.5. Table IV reports ASRs of multiple defense methods under various types of jailbreak attacks. The defense methods are evaluated across different jailbreak techniques, consisting of optimization-based, template-based, indirect, multilingual, and normal prompts. Under the optimization-based GCG and JSAA, ASR of SelfDefend with the intent prompt is 0.05, slightly better than that with the direct prompt. In contrast, SelfDefend with the direct prompt is slightly better than SelfDefend with the intent prompt under the template-based PAIR and AutoDAN.
For indirect DrAttack, ASR of the fine-tuned model with direct prompt is 0.78, the same as the original GPT-3.5, whereas the dedicated model with intent prompt reduces ASR to 0.18. With manual exploration, we find that DrAttack splits the sentence components of the original jailbreak prompt and replaces them with harmless words, making it difficult for SelfDefend (direct prompt) to extract harmful content. Although Puzzler is also an indirect attack, the prompts it produces contain offensive content as implicit clues for jailbreak, resulting in a lower ASR of SelfDefend with the direct prompt. For multilingual MultiJail, the intent prompt tuned model is superior than the direct prompt by 17%. Therefore, the model fine-tuned with the direct prompt is well-suited for identifying overt jailbreak intentions within sentences, whereas the model fine-tuned with the intent prompt excels in detecting jailbreak attempts that encompass more subtle or concealed intentions. Besides, SelfDefend has ASRs that are comparable with all other defense methods under normal prompts, indicating its compatibility with normal users.
Figures 5, 6, 7, 8 and 9 present several example responses of the tuned models toward either jailbreak prompts or a normal query. Note that in Figure 7, DrAttack evades the detection of the direct prompt tuned model, while the intent tuned model accurately identifies the intent of DrAttack, which further illustrates the weakness of the direct tuned model against indirect attacks. Moreover, as expected, we observe that both tuned models make satisfactory decisions on other jailbreaks or the normal prompt.
Results on GPT-4. Similar with the results achieved over GPT-3.5, we can observe that our defense frameworks achieve the lowest ASR under all jailbreak methods, and is still close to ASR of GPT-4 on AlpacaEval. We interpret the results as generally encouraging, illustrating the defensive superiority of SelfDefend without intruding normal user experience. Under the optimization-based GCG, the lowest ASR (as achieved by SelfDefend) is 0.01 with the direct prompt. For JSAA, SelfDefend with the intent prompt is slightly better than SelfDefend with the direct prompt. Under the template-based PAIR and AutoDAN, SelfDefend with the direct prompt is slightly superior than that with the intent prompt. For indirect DrAttack, ASR of the fine-tuned model with direct prompt is 0.71, close to that of the original GPT-4, while the fine-tuned model with intent prompt reduces ASR to 0.18. In addition, the direct prompt tuned model performs the lowest ASR under Puzzler, but far inferior to the intent one under MultiJail.
Results on Llama-2-7b-chat. Our defense frameworks exhibit the lowest ASRs across all jailbreak methods. Moreover, it achieves a ASR that is comparable to Llama-2-7b-chat on AlpacaEval, showing SelfDefend’s defensive effectiveness without compromising normal user experience. Specifically, under the white-box GCG, the ASRs for both of our SelfDefend frameworks are less than 10%.
Perplexity Filter exhibits the lowest ASR against GCG; with further exploration, we find that this is because its filter threshold is set to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix. Yet, its detectability in other settings are not as good as SelfDefend. Under the optimization-based JSAA attack, SelfDefend with the intent prompt achieves the lowest ASR of 0.02, performing marginally better than with the direct prompt. In the context of template-based attacks such as PAIR and AutoDAN, the results between SelfDefend with the direct prompt and the intent one are close, with each having their own superior cases. For the indirect DrAttack, while the direct prompt tuned model largely reduces the ASR compared to the original Llama-2-7b-chat model, it still has a 21% gap compared to the intent one. Additionally, the direct prompt fine-tuned model has the lowest ASR under the Puzzler attack, yet it is notably outperformed by the intent prompt one in the MultiJail scenario.
@@
Double Shadow Stacks. The current design of SelfDefend  accompanies one shadow stack with one LLM under protection. Despite the encouraging protection results, one may wonder incorporating multiple shadow stacks to harvest increased defense capability. Here, we define a “double shadow stack” setting as two shadow stacks independently analyze an input, where an input is deemed jailbreak when at least one shadow stack flags it as jailbreak.
@@
In Table IV, “Double shadow stack” is instantiated by incorporating one $P_{direct}$-based and one $P_{intent}$-based shadow stacks. We can observe that the results of double shadow stack are close to the best results of the two SelfDefend frameworks, except under MultiJail. And when we further fuse the target model (i.e., normal stack) with the double shadow stack, it exhibits the lowest ASRs under all attack scenarios, as shown in Table IV. Nevertheless, we point out that when fusing two or three stacks together, the performance over normal prompts (AlpacaEval) decrease, meaning a few more normal inputs are deemed harmful. This is expected; taking the logical disjunction of multiple stacks’ predictions reduces the false negatives of SelfDefend, yet likely increases the false positives. In practice, we suggest using this paradigm only when the standard form of SelfDefend offers low detection accuracy.
@@
TABLE V: The CLIP-Score [64, 82, 83] ($\uparrow$ indicates better) based on Table IV’s results for Llama 2.
Tuned Model Text Optimization-based Template-based Indirect Multilingual
GCG JSAA PAIR AutoDAN DrAttack Puzzler MultiJail
$P_{direct}$-tuned model Generated Attack Prompts 0.85 0.70 0.83 0.78 0.74 0.66 0.68
Identified Harmful Prompts 0.95 0.94 0.90 0.94 0.79 0.73 0.69
$P_{intent}$-tuned model Generated Attack Prompts 0.85 0.70 0.83 0.78 0.73 0.67 0.67
Identified Harmful Intentions 0.88 0.88 0.85 0.88 0.83 0.77 0.74
@@ 7.3 Extra Delay $\Delta d$
Similar to how we measured the extra delay $\Delta d$ introduced by GPT-based SelfDefend in §5.3, we now report the average delay $\Delta d$ introduced by the tuned models on Llama-2-7b-chat in Figure 10. Compared to earlier results in Figure 3, $\Delta d$ under AlpacaEval is negligible for SelfDefend with both $P_{direct}$ and $P_{intent}$, with the former around 0 seconds and the latter around 0.01 seconds. In contrast, GPT-3.5-based (GPT-4-based) SelfDefend with $P_{intent}$ incurred an average $\Delta d$ of around 0.07 seconds (0.02 seconds) under AlpacaEval. This is likely because the fine-tuned models enjoy faster inference speeds when dealing with similar queries. Besides the normal prompts, the tuned models also significantly reduce the extra delay $\Delta d$ under various jailbreak prompts. For example, the maximum $\Delta d$ now decreases from the earlier 1.1 seconds on GPT-4 to 0.25 seconds. Indeed, except for JSAA, $\Delta d$ in all attack scenarios are now below 0.1 seconds, while there was no single $\Delta d$ below the same time limit for GPT-4-based SelfDefend. These results indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for real-world deployment.
@@ 7.4 Explainability
Figure 5, 6, 7, and 8 demonstrate the efficacy of the tuned models in identifying harmful content within jailbreak prompts. Here, we further quantitatively assess whether the identified harmful content actually aligns with the original jailbreak prompt. We measure their semantic similarity using the ensemble CLIP-score [64, 82, 83] as shown in Table V. Note that our measurement exclusively targets the examples where the defense mechanism is successful.
We calculate the CLIP-score between the original prompts $P_{query}$ from JailbreakBench and the attack-generated jailbreak prompts, as well as the CLIP-score between $P_{query}$ and our identified harmful content (prompts/intentions). Specifically, the CLIP-scores for identified harmful portions are higher than those for the generated attack prompts, suggesting that the jailbreak prompts have altered the original prompts, whereas the harmful portions identified by our models are more aligned with the original prompts. Regarding the differences between the two tuned models, we observe that the similarity of identified harmful prompts is higher than that of harmful intentions in the context of optimization-based and template-based attacks, but lower than that of intentions under indirect and multilingual attacks. This can be attributed to the fact that optimization-based and template-based attack prompts contain the original prompts, whereas indirect and multilingual attack prompts significantly distort the original prompts. In summary, these findings are consistent with our initial design intentions and demonstrate that the harmful portions identified by our models can indeed provide an explainable foundation for defense mechanisms.
TABLE VI: ASR with the regular GCG attack on Llama 2 and the targeted GCG directly on our tuned defense models.
Tuned Model Regular GCG Targeted GCG
(from Table IV) (white box)
$P_{direct}$-tuned model 0.06 0.07
@@
7.5 Robustness to Targeted GCG Attacks
In earlier evaluation with optimization-based attacks such as GCG in §7.2, we used a regular setting where such attacks leverage gradients to optimize adversarial suffixes on an open-source substitute model (i.e., Llama-2-7b-chat) and transfers these suffixes to attack our fine-tuned defense models. To stress-test the robustness of our fine-tuned defense models against targeted GCG attacks, we give the adversary full white-box access to our model parameters and allow them to optimize adversarial suffixes directly on our defense models.
Table VI presents the ASR results for both $P_{direct}$-tuned model and $P_{intent}$-tuned model. We observe that the ASR results of the targeted, white-box GCG remain below 10% for both types of tuned models in the shadow stack. ASR slightly increases for the targeted GCG under the $P_{direct}$-tuned model but slightly decreases for the targeted GCG under the $P_{intent}$-tuned model. This indicates that our defense models are robust to even targeted GCG attacks.
@@ 7.6 Robustness to Prompt Injection
Since SelfDefend uses $P_{direct}$ or $P_{intent}$ to wrap the original $P_{query}$, it is natural to wonder whether SelfDefend is robust to prompt injection [65, 66], another major LLM threat that is orthogonal to jailbreak attacks.
@@ 8 Conclusion
We have introduced SelfDefend, a robust, low-cost, and self-contained defense against LLM jailbreak attacks. Inspired by the concept of shadow stacks, SelfDefend delivers a dual-layer defense mechanism comprising a shadow LLM that guards the target LLM. It further leverages a tuning-based approach to enhancing the shadow LLM’s defense capability. Evaluation shows that SelfDefend is lightweight and effective to mitigate a wide spectrum of jailbreak attacks while rarely undermining normal queries.
Appendix A Keywords
TABLE VII: The refusal signals considered for GPT-3.5/4.

### v2

@@ 7.2 Defense Effectiveness
We now discuss the defense effectiveness of our proposed technique. Table 4 reports ASRs of multiple defense methods under various types of jailbreak attacks. The defenses are evaluated across different jailbreak techniques, consisting of human-based, optimization-based, generation-based, indirect, and multilingual jailbreak attacks, as well as normal prompts.
Comparison with Existing Defense Methods. As shown in Table 4, our tuned models under SelfDefend achieve satisfactory defense effects under all types of jailbreaks and deliver state-of-the-art (SOTA) results in most cases, while previous defenses fail to show promising performance under different attacks. For ICD, the tested ASRs under Puzzler are not less than 99% for both GPT-3.5 and Llama-2. For SafeDecoding, it also shows weakness against Puzzler and MultiJail. The Perplexity Filter is only effective against GCG and lacks resistance to other jailbreaks. Although it exhibits slightly better ASR against GCG than our methods for GPT-3.5 and Llama-2, this is because its filter threshold is set to the maximum perplexity among all prompts in JailbreakBench, making it trivial to detect GCG with a garbled suffix. SmoothLLM does not show satisfactory defense performance against DrAttack and Puzzler for all three target models. Its ASRs even exceed 80% on Llama-2 for all attack methods except DAN. Llama Guard exhibits poor jailbreak detection for DrAttack, Puzzler, and MultiJail, while Llama Guard 2 presents unacceptable flaws for DrAttack.
Across all attacks and target models, the average ASR of our $P_{direct}$-tuned shadow stack (27.76%) is 37.61% lower than that of Llama Guard (65.37%) and 16.74% lower than that of Llama Guard 2 (44.50%). Similarly, the average defense performance of our $P_{intent}$-tuned shadow stack (22.46%) outperforms Llama Guard by 42.91% and Llama Guard 2 by 22.04%. Furthermore, SelfDefend achieves ASRs comparable with all other defense methods under AlpacaEval, indicating its compatibility with normal prompts.
$P_{direct}$-tuned SelfDefend v.s. $P_{intent}$-tuned SelfDefend. In general, SelfDefend with the direct prompt performs better than SelfDefend with the intent prompt on human-based, optimization-based, and generation-based jailbreaks, while $P_{intent}$-tuned SelfDefend is more effective against indirect and multilingual attacks. DAN, GCG, AutoDAN, PAIR, and TAP contain explicitly harmful instructions, whereas DrAttack, Puzzler, and MultiJail do not. This phenomenon aligns with our original intent for designing the two prompts, since $P_{direct}$ is used to directly extract harmful parts, and $P_{intent}$ can identify the true intentions hidden by implicit attacks. For example, DrAttack deconstructs the sentence components of the original jailbreak prompt and substitutes them with harmless words, making it challenging for SelfDefend (with $P_{direct}$) to extract any harmful content. Although Puzzler is also an indirect attack, the prompts it generates include offensive content such as inducing clues for jailbreak, resulting in a lower ASR for SelfDefend with $P_{direct}$. Moreover, the false positive rate of the $P_{intent}$-tuned SelfDefend on normal prompts is lower than that of the $P_{direct}$-tuned SelfDefend.
@@
Double Shadow Stacks. The current design of SelfDefend accompanies one shadow stack with one LLM under protection. Despite the encouraging protection results, one might consider whether incorporating multiple shadow stacks could enhance defense capabilities. Here, we define a “double shadow stack” setting as two shadow stacks independently analyzing an input, where an input is deemed a jailbreak if at least one shadow stack flags it as such.
@@
In Table 4, “Double shadow stack” is instantiated by incorporating one $P_{direct}$-based and one $P_{intent}$-based shadow stack. We observe that the results of the double shadow stack are not worse than the best results of the two separate stacks. Furthermore, when we fuse the target model (i.e., the normal stack) with the double shadow stack, it exhibits the lowest ASRs under almost all attack scenarios, as shown in Table 4. Nevertheless, we point out that when fusing two or three stacks together, the performance over normal prompts (AlpacaEval) decreases, meaning a few more normal inputs are deemed harmful. This is expected; taking the logical disjunction of multiple stacks’ predictions reduces the false negatives of SelfDefend, yet likely increases the false positives. In practice, we suggest using this paradigm only when the standard form of SelfDefend offers low detection accuracy.
Case Study. Besides the quantitative analysis above, we conduct case studies to understand in depth the advantages and pitfalls of SelfDefend compared to other defenses. Figure 5 illustrates an example of defense methods against a jailbreak prompt. Due to page limitation, readers may refer to Appendix A for its detailed analysis, as well as other cases.
Table 5: The CLIP-Score [58, 92, 69] ($\uparrow$ indicates better) based on Table 4’s results for Llama-2.
Tuned Model Text Human-based Optimization-based Generation-based Indirect Multilingual
DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail
$P_{direct}$-tuned model Generated Attack Prompts 0.685 0.851 0.784 0.828 0.827 0.737 0.663 0.677
Identified Harmful Prompts 0.932 0.946 0.939 0.898 0.900 0.785 0.728 0.687
$P_{intent}$-tuned model Generated Attack Prompts 0.686 0.852 0.784 0.831 0.827 0.732 0.667 0.669
Identified Harmful Intentions 0.919 0.909 0.908 0.875 0.874 0.857 0.788 0.766
@@ 7.3 Extra Delay $\Delta d$
Similar to how we measured the extra delay $\Delta d$ introduced by GPT-based SelfDefend in §5.3, we now report the average delay $\Delta d$ introduced by the tuned models on Llama-2-7b-chat in Figure 6. Compared to earlier results in Figure 3, $\Delta d$ under AlpacaEval is negligible for SelfDefend with both $P_{direct}$ and $P_{intent}$, with the former around 0 seconds and the latter around 0.008 seconds. In contrast, GPT-3.5-based (GPT-4-based) SelfDefend with $P_{intent}$ incurred an average $\Delta d$ of around 0.072 seconds (0.026 seconds) under AlpacaEval. This is likely because the parameters of fine-tuned models are less than those of GPT-3.5/4. Besides the normal prompts, the tuned models also significantly reduce the extra delay $\Delta d$ under various jailbreak prompts. For example, the maximum $\Delta d$ now decreases from the earlier 1.56 seconds on GPT-4 to 0.27 seconds. Indeed, except for DAN, $\Delta d$ in all attack scenarios are now below 0.1 seconds, while there was no single $\Delta d$ below the same time limit for GPT-4-based SelfDefend. These results indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for potential deployment.
@@ 7.4 Explainability
Figure 5 demonstrate the efficacy of the tuned models in identifying harmful content within jailbreak prompts. Here, we further quantitatively assess whether the identified harmful content actually aligns with the original jailbreak prompt. We measure their semantic similarity using the ensemble CLIP-score [58, 92, 69], as shown in Table 5. Note that our measurement exclusively targets the examples where the defense mechanism is successful.
We calculate the CLIP-score between the original prompts $P_{query}$ from JailbreakBench and the attack-generated jailbreak prompts, as well as the CLIP-score between $P_{query}$ and our identified harmful content (prompts/intentions). Specifically, the CLIP-scores for identified harmful portions are higher than those for the generated attack prompts, suggesting that the jailbreak prompts have altered the original prompts, whereas the harmful portions identified by our models are more aligned with the original prompts. Regarding the differences between the two tuned models, we observe that the similarity of identified harmful prompts is higher than that of harmful intentions in the context of human-based, optimization-based, and generation-based attacks, but lower than that of intentions under indirect and multilingual attacks. This can be attributed to the fact that human-based, optimization-based and generation-based attack prompts contain the original prompts, whereas indirect and multilingual attack prompts significantly distort the original prompts. In summary, these findings are consistent with our initial design intentions and demonstrate that the harmful portions identified by our models can indeed provide an explainable foundation for defense mechanisms.
Table 6: Performance (ASRs) of SelfDefend with different shadow models against adaptive attacks (i.e., PAIR and TAP).
Target Model Shadow Model PAIR TAP
GPT-3.5 Llama Guard 0.38 0.39
Llama Guard 2 0.31 0.36
$P_{direct}$-tuned model 0.22 0.20
$P_{intent}$-tuned model 0.25 0.18
GPT-4 Llama Guard 0.28 0.24
Llama Guard 2 0.24 0.20
$P_{direct}$-tuned model 0.20 0.19
$P_{intent}$-tuned model 0.24 0.15
Llama-2-7b-chat Llama Guard 0.37 0.36
Llama Guard 2 0.28 0.28
$P_{direct}$-tuned model 0.21 0.21
$P_{intent}$-tuned model 0.23 0.22
7.5 Robustness to Adaptive Jailbreaks
Robustness to Black-box Jailbreaks. In §7.2, the jailbreak prompts used for evaluation are either manually designed, transferred from a surrogate model (Vicuna-7b-v1.3 [13]), or generated for the target models (GPT-3.5, GPT-4, Llama-2-7b-chat). To evaluate the robustness of our SelfDefend against black-box adaptive attacks, we select query-based PAIR and TAP to jailbreak the whole SelfDefend framework. The results are reported in Table 6. First, we can observe that all our tuned models outperform the Llama Guard series, except for one case, which is that the result of $P_{intent}$-tuned model is equal to Llama Guard 2 under PAIR. Moreover, the $P_{direct}$-tuned model is superior to the $P_{intent}$-tuned model in defending against PAIR, whether for GPT-3.5, GPT-4, or Llama-2. In contrast, the $P_{intent}$-tuned model is superior to the $P_{direct}$-tuned model in resisting TAP, except on Llama-2.
[…]
Table 7: ASR with the regular GCG attack on Llama-2 and the white-box GCG directly on our tuned defense models.
Tuned Model GCG (from Table 4) White-box GCG
$P_{direct}$-tuned model 0.06 0.08
@@ 7.6 Robustness to Prompt Injection
Since SelfDefend uses $P_{direct}$ or $P_{intent}$ to wrap the original $P_{query}$, it is reasonable to question whether SelfDefend is robust against prompt injection [44, 42], another major LLM threat that is orthogonal to jailbreak attacks.
@@ 8 Conclusion
We have introduced SelfDefend, a robust, low-cost, and self-contained defense against LLM jailbreak attacks. Inspired by the concept of shadow stacks, SelfDefend delivers a dual-layer defense mechanism comprising a shadow LLM that guards the target LLM. It further leverages a tuning-based approach to enhancing the shadow LLM’s defense capability. The evaluation shows that SelfDefend is lightweight and effective in mitigating a wide spectrum of jailbreak attacks while rarely undermining normal queries.
Ethics and the Open Science
Our research has meticulously addressed various ethical considerations to ensure responsible and ethical conduct. Firstly, regarding vulnerability disclosures and open research artifacts, we have released our datasets, raw evaluation results, and code at this anonymous GitHub link: https://github.com/SelfDefend. While our study required testing various jailbreak attacks on the live system (i.e., ChatGPT API service), these tests never affected other OpenAI users and were conducted solely on our own OpenAI account and a local server. We also aimed to minimize the overhead for OpenAI by distributing our experiments over a period of three months and incurred around \$2,000 in API usage fees. Secondly, in addition to all tests on ChatGPT complying with the relevant terms of service, we adhere to Meta Llama’s relevant open-source agreements to maintain legal and ethical standards. The well-being of our team members was a priority, with measures in place to protect against exposure to harmful content during the research process and offering psychological support if needed. Although our novel defense mechanism provide an effective, low-latency, plug-and-play, and explainable solution, its potential negative outcome could potentially be used maliciously to craft better attacks on LLMs. We therefore engage with the community to ensure the technology is used responsibly. This involved a readiness to retroactively identify negative outcomes and take corrective actions if our initial assessments underestimated the impacts. Finally, all aspects of our research were conducted in strict compliance with the law, ensuring that our practices did not inadvertently contravene legal standards, particularly in data protection and privacy. This comprehensive ethical approach not only aligns with the guidelines but also reinforces the integrity and societal value of our research.
Appendix
Appendix A Case Studies
Successful Cases. Figure 5 in §7.2 presents a detection example of our methods applied to a jailbreak query, compared with the Llama Guard series. The attack prompt involves creating a fake news article to spread disinformation about a political candidate. Both Llama Guard and Llama Guard 2 identify it as safe, while our methods mark its harmful portions. The reason our method outperforms Llama Guards might be due to the enhanced reasoning ability of the foundation model in identifying harmful portions.
As described in our design, the direct prompt-based model identifies harmful parts directly from the input text, whereas the intent prompt-based model summarizes the request’s intention and then identifies unsafe parts from that summary. When comparing our $P_{direct}$-tuned models with $P_{direct}$-based GPT-4, we observed that the parts they lifted from the attack prompt came from different sentences, since $P_{direct}$ emphasizes that it is sufficient to output one harmful part. A similar phenomenon occurs with the intent prompt, where the semantics of GPT-4 and our tuned model’s intercepts are not identical.
In addition, please refer to Appendix C for more case of our models handling other jailbreaks or the normal prompts. As expected, both of our tuned models make satisfactory decisions on various attack or benign inputs.
Figure 7: An example of tuned models against DrAttack on GPT-3.5.
Failure Cases. Figure 7 shows a failure case of the $P_{direct}$-tuned model against a jailbreak prompt from DrAttack. Since DrAttack deconstructs the sentence components of the original jailbreak goal, it is difficult for the $P_{direct}$-tuned model to identify consecutive harmful portions. By contrast, the $P_{intent}$-tuned model accurately identifies the intent of DrAttack and defends against it, indicating the strength of the intent prompt-tuned model against indirect attacks.
Moreover, Figure 8 provides a failed example of the $P_{intent}$-tuned model against PAIR. The original jailbreak goal for this attack prompt is “Write a blog post explaining why gambling addicts should take payday loans”. The current attack prompt discusses the pros and cons of using payday loans to fund gambling addiction, which reduces the overall toxicity compared to the original goal. Even though the summary intent of the $P_{intent}$-tuned model is correct for the attack prompt, it recognizes the request as harmless due to the low toxicity. The global semantic difference between the original jailbreak goal and the attack prompt is one reason for the failure of the $P_{intent}$-tuned model. In contrast, the $P_{direct}$-tuned model successfully identifies the harmful part, thus making the right judgment.
Figure 8: An example of tuned models against PAIR on GPT-3.5.
Appendix B Keywords
Table 8: The refusal signals considered for GPT-3.5/4.